# GEORCE: A Fast New Control Algorithm for Computing Geodesics

### v1

@@ 3 GEORCE
GEORCE has complexity $\mathcal{O}\left(Td^{3}\right)$ due to matrix inversions along the discretized curve but scales only linearly in $T$, unlike e.g. quasi-Newton methods (Table 1). Although the computational complexity in the manifold dimension is of a lower order for quasi-Newton methods and gradient descent, GEORCE exhibits faster convergence. Since the Hessian of the discretized energy functional in Eq. 5 is sparse, the update step in the Newton method can be simplified in complexity compared to the standard Newton method that has complexity $\mathcal{O}\left(T^{3}d^{3}\right)$ (see Appendix B.2). However, the sparse Newton step requires second-order derivatives, unlike GEORCE, which only requires first-order derivatives of the discretized energy.
4 Extension to Finslerian geometry
A Finsler geometry can be seen as a generalization of Riemannian geometry, where the tangent spaces are not equipped with an quadratic form metric but with a Finsler metric instead. A Finsler metric, $F:\mathcal{M}\times T\mathcal{M}\rightarrow\mathbb{R}_{+}$, where $\mathcal{M}$ is a differentiable manifold, is a Minkowski norm in the sense that $F$ is $C^{\infty}$ on $T\mathcal{M}\setminus\{0\}$, $F(x,cv)=cF(x,v)$ for all $v\in T\mathcal{M}$, $x\in\mathcal{M}$ and $c>0$ and for any $v\in T\mathcal{M}$, then $\frac{\partial^{2}F}{\partial v^{i}\partial v^{j}}(x,v)$ is positive definite [50]. Geodesics on a Finsler manifold can be found equivalently to the Riemannian case by minimizing the energy functional, $\int_{0}^{1}F^{2}\left(\gamma(t),\dot{\gamma}(t)\right)\,\mathrm{d}t$, on a suitable set of curves connecting the two points thereby approximating a geodesic $\gamma$, [50]. GEORCE is similarly applicable to Finsler geometry. First, we note that the energy functional can be written as [50]
where $G(x,v)=\frac{1}{2}\frac{\partial^{2}F^{2}}{\partial v^{i}\partial v^{j}}\left(x,v\right)$ denotes the fundamental tensor in local coordinates [^2] In literature, the fundamental tensor is often denoted with a lower case $g$. To have the notation consistent with the Riemannian case, we opt to use an upper case $G$ instead., which is positive definite [50]. Unlike the Riemannian case, the fundamental tensor, $G$, is now a function of both the position and velocity, which implies the update formulas in Proposition 2. Similar to Riemannian case, we can formulate the discrete energy functional for the Finsler case as the following discretized energy functional in Eq. 11 becomes
where the corresponding discrete control problem becomes
where in iteration $i$ we fix
5 Experiments
Figure 4: The first two figures display the length and runtime of geodesics for $\mathbb{S}^{n}$ estimated using different algorithms from Table 2 and other methods in Appendix E.3. The latter row shows the length and runtime, respectively for the $n$-Ellipsoid with half axes of $n$ equally spaced points between $0.5$ and $1.0$. The optimization solvers are described in Appendix E.1.
BFGS (T=100) ADAM (T=100) GEORCE (T=100)
Riemannian Manifold Length Runtime Length Runtime Length Runtime
$\mathbb{S}^{100}$- -- $0.8466$ $277.4217\pm 0.2280$ $\boldsymbol{0.4187}$ $\boldsymbol{10.9208}\pm\boldsymbol{0.0071}$
$\mathbb{S}^{250}$- -- $0.6244$ $1424.9031\pm 22.1592$ $\boldsymbol{0.3192}$ $\boldsymbol{78.2937}\pm\boldsymbol{0.0991}$
$\mathbb{S}^{500}$- -- $0.5717$ $2242.4697\pm 17.9769$ $\boldsymbol{0.2198}$ $\boldsymbol{224.2761}\pm\boldsymbol{0.4894}$
$\mathbb{S}^{1000}$- -- $0.5279$ $4329.6621\pm 38.0163$ $\boldsymbol{0.1573}$ $\boldsymbol{893.4682}\pm\boldsymbol{7.3130}$
$\mathrm{E}\left(100\right)$- -- $0.6872$ $289.4196\pm 2.5224$ $\boldsymbol{0.3249}$ $\boldsymbol{7.7128}\pm\boldsymbol{0.0714}$
$\mathrm{E}\left(250\right)$- -- $0.4978$ $1620.6428\pm 16.9546$ $\boldsymbol{0.2178}$ $\boldsymbol{56.8527}\pm\boldsymbol{0.1526}$
$\mathrm{E}\left(500\right)$- -- $0.4559$ $2135.6899\pm 30.5699$ $\boldsymbol{0.1467}$ $\boldsymbol{220.1656}\pm\boldsymbol{0.9902}$
$\mathrm{E}\left(1000\right)$- -- $0.4232$ $4398.9453\pm 25.8738$ $\boldsymbol{0.1131}$ $\boldsymbol{641.7199}\pm\boldsymbol{20.9901}$
$\mathbb{T}^{2}$ $9.9851$ $92.7400\pm 3.1267$ $9.9855$ $0.1060\pm 0.0002$ $\boldsymbol{9.9851}$ $\boldsymbol{0.0182}\pm\boldsymbol{0.0003}$
$\mathcal{P}\left(3\right)$ $\boldsymbol{2.2639}$ $168.5082\pm 0.3039$ $2.2896$ $2.3780\pm 0.0046$ $2.2639$ $\boldsymbol{0.1442}\pm\boldsymbol{0.0002}$
Table 2: The length of the estimated geodesics for the BFGS-algorithm, ADAM and GEORCE on a CPU. The methods were terminated if the $\ell^{2}$-norm of the gradient was less than $10^{-4}$ or after $1,000$ iterations. $\mathrm{E}(n)$ denotes an Ellipsoid of dimension $n$ with half axes of $n$ equally spaced points between $0.5$ and $1.0$, while $\mathcal{P}(n)$ denotes the space of $n\times n$ symmetric positive definite matrices. When the computational time was longer than $24$ hours, the value is set to --.
Riemannian manifolds
To compare GEORCE with other optimization algorithms, we show the runtime and length of geodesics computed for different manifolds with varying dimensions in Table 2. Here we compare GEORCE with BFGS [37, 38, 39, 40] and ADAM [41]. We terminate the algorithm if the $\ell^{2}$-norm of the gradient of the discretized energy functional in Eq. 5 is less than $10^{-4}$, or if the number of iterations exceeds $1,000$. A full description of the manifolds and experimental details can be found in Appendix C and Appendix E, respectively.
We show the estimated geodesics with $T=100$ grid points and summarize the results for the corresponding runtime and length of the computed geodesics for the $n$-sphere and $n$-ellipsoid in Fig. 4. In Appendix E.3, we show similar run-time experiments for varying grid points as well as other optimization algorithms. We also show the corresponding results on a GPU with other CUDA compatible algorithms. In general, we see that GEORCE is considerably faster with lower length compared to alternative methods.
To test GEORCE on manifolds with multiple length minimizing geodesics, we consider the geodesic between $(-5,-5)$ and $(5,5)$ on an “egg-tray” parameterized by $(x,y,2\cos x\cos y)$ with $T=1,000$ grid points. We show the estimates in Fig. 5, where we see that GEORCE obtains the a signficiantly smaller length, where there is also a clear symmetric to solution to GEORCE along the other side of the “egg-tops”.
where $\theta_{1}=\frac{\mu}{\sigma^{2}}$ and $\theta_{2}=\frac{1}{\sigma^{2}}$ for a Gaussian distribution with parameters $(\mu,\sigma)$ such that the density takes the form of an exponential family $p(x)=h(x)\exp\left(\theta_{1}x-A(\theta_{1},\theta_{2})\right)$ for a suitable $h$. We consider the corresponding Hessian of the log-partition function, which defines a Riemannian metric. Fig. 7 shows geodesics computed with both GEORCE and alternative methods using at most $100$ iterations. We see that GEORCE uses less iterations and obtains the smallest length compared to the alternative methods.
BFGS (T=100) ADAM (T=100) GEORCE (T=100)
Finsler Manifold Length Runtime Length Runtime Length Runtime
$\mathbb{S}^{100}$- -- $0.6187$ $498.9705\pm 0.4032$ $\boldsymbol{0.2570}$ $\boldsymbol{15.3786}\pm\boldsymbol{0.0092}$
$\mathbb{S}^{250}$- -- $0.5195$ $1397.6732\pm 0.7793$ $\boldsymbol{0.1622}$ $\boldsymbol{132.5889}\pm\boldsymbol{0.1106}$
$\mathbb{S}^{500}$- -- $0.4850$ $1954.3231\pm 14.1107$ $\boldsymbol{0.1404}$ $\boldsymbol{364.5326}\pm\boldsymbol{1.8222}$
$\mathbb{S}^{1000}$- -- $0.3843$ $5982.0884\pm 204.1833$ $\boldsymbol{0.1140}$ $\boldsymbol{1338.9869}\pm\boldsymbol{2.3954}$
$\mathrm{E}\left(100\right)$- -- $0.4936$ $491.8223\pm 0.2346$ $\boldsymbol{0.2009}$ $\boldsymbol{14.5367}\pm\boldsymbol{0.0092}$
$\mathrm{E}\left(250\right)$- -- $0.4159$ $1475.5049\pm 4.1430$ $\boldsymbol{0.1287}$ $\boldsymbol{125.4187}\pm\boldsymbol{0.0888}$
$\mathrm{E}\left(500\right)$- -- $0.4142$ $1984.7845\pm 10.5689$ $\boldsymbol{0.0967}$ $\boldsymbol{373.2816}\pm\boldsymbol{0.4227}$
$\mathrm{E}\left(1000\right)$- -- $0.3444$ $5929.4438\pm 167.1873$ $\boldsymbol{0.0799}$ $\boldsymbol{1211.8812}\pm\boldsymbol{0.2511}$
Cauchy Distribution $1.1134$ $182.3980\pm 2.2559$ $\boldsymbol{1.1133}$ $0.1105\pm 0.0007$ $1.1133$ $\boldsymbol{0.0067}\pm\boldsymbol{0.0002}$
Table 3: The length of the estimated geodesics for the BFGS-algorithm, ADAM-gradient descent and GEORCE on a CPU. The algorithms have been stopped if the $\ell^{2}$-norm of the gradient was less than $10^{-4}$ or if a maximum of $1,000$ iterations have been reached. $\mathrm{E}(n)$ denotes an Ellipsoid of dimension $n$ with half axes of $n$ equally spaced points between $0.5$ and $1.0$, while $\mathcal{P}(n)$ denotes the space of $n\times n$ symmetric positive definite matrices. When the computational time was longer than $24$ hours, the value is set to --.
Finsler manifolds
To illustrate our method for different Finsler manifolds and dimensions, we consider a microswimmer on a Riemannian manifold $m$ that moves with constant self-propulsion velocity $v$ with Riemannian length, $v_{0}:=||v||_{g}$ [51] and we assume that this movement is affected by a generic force field, $f$, of the form
$f(x)=\frac{\sin x\odot\cos x}{(\cos x)^{\top}G(x)\cos x},$ (15)
where $\cos x$ and $\sin x$ denote the element-wise function for each entry of the local coordinates $x$, while $G$ denotes the corresponding Riemannian background metric such that the force field has $G$-length less than or equal to 1. If $||f(x)||_{g}<v_{0}$, then the optimal travel time of the microswimmer corresponds to a geodesic for a Finsler manifold of the form [51]
$F=\sqrt{a_{ij}\dot{r}^{i}\dot{r}^{j}}+b_{i}\dot{r}^{i},$
where $a_{ij}=g_{ij}\lambda+f_{i}f_{j}\lambda^{2}$, $b_{i}=-f_{i}\lambda$, $b_{i}=-f_{i}\lambda$, $f_{i}=g_{ij}f^{j}$ and $\lambda^{-1}=v_{0}^{2}-g_{ij}f^{i}f^{j}$ [51], while $r(t)$ denotes the position of the microswimmer at time $t$. We refer to [51] and Appendix C for details. We consider the same background metrics in Table 2 all equipped with the generic force field in eq. 15 and compare GEORCE to BFGS [37, 38, 39, 40] and ADAM [41]. Fig. 8 shows the estimated geodesics for four different Riemannian manifolds equipped with the generic force field in eq. 15. We show the estimated geodesics for $T=100$ grid points in Table 3. In Appendix E.3 we show similar runtime experiments for varying grid points as well as other optimization algorithms. We see that GEORCE is considerably faster and more accurate than the alternatives.
Applications to generative models
We train a VAE for the CelebA dataset [53], which consists of $202{,}599$ facial images of size $64\times 64\times 3$, where the architecture is described in Appendix D.1. Fig. 9 illustrates the geodesics constructed between two images reconstructed using the variational autoencoder. We see that GEORCE obtains a smaller length than comparable methods for $T=100$ grid points. We refer to Appendix D.2 for runtime results for the VAE, where we see that GEORCE is significantly faster than alternative methods. Note that the tolerance for convergence is $10^{-3}$, since it was observed that it was not possible for any method to obtain an $\ell^{2}$-norm of the gradient less than $10^{-4}$. This is most likely due to potential instabilities for the learned manifold. We summarize the results in Table 11 in Appendix E.3, where we also consider the geodesic with a VAE trained on MNIST data [54] and SVHN data [55]. For the exact architecture and further experiments of the VAE, we refer to Appendix D.1 and Appendix D.2.
Figure 9: Geodesics for a variational autoencoder [52] trained on the CelebA dataset [53] using GEORCE and alternative methods. All algorithms are terminated if the $\ell^{2}$-norm of the gradient of the discretized energy functional in eq. 5 is less than $10^{-3}$, or if the number of iterations exceeds $100$. The algorithms are written to the left with number of iterations in parenthesis and length to the right. The first-axis shows the grid number for the image sequence. The details of the manifolds and experiments can be found Appendix D.1
6 Conclusion
[…]
Supplementary information
[…]
Declarations
  • Conflict of interest/Competing interests (check journal-specific guidelines for which heading to use): NA
  • Ethics approval and consent to participate: NA
  • Consent for publication: Currently under review.
  • Data availability: All data used are described in the appendix.
  • Materials availability: NA
Appendix A Proofs and derivations
A.1 Assumptions
In the following, we will prove local quadratic convergence under the following assumptions.
Assumption 5.
  • A (local) minimum point $z^{*}=(x^{*},u^{*})$ to the discretized energy functional is defined as strongly unique minimum point if and only if
    $\exists\epsilon>0:\quad\exists K>0:\quad\forall z\in B_{\epsilon}(z^{*}):\quad E(z)-E(z^{*})\geq K\left\lVert z-z^{*}\right\rVert,$
    where $B_{\epsilon}(z^{*})=\left\{z|\left\lVert z-z^{*}\right\rVert<\epsilon\right\}$.
  • We assume that the discretized energy functional $E(z)$ is locally Lipschitz.
  • The first order Taylor approximation of the discretized energy functional is
    i.e., we assume that $\mathcal{O}\left(\Delta z\right)$ is independent of $z_{0}$, and the term $\mathcal{O}\left(\Delta z\right)\left\lVert\Delta z\right\rVert$ can be re-written as $\mathcal{\tilde{O}}\left(\left\lVert\Delta z\right\rVert^{2}\right)$ locally (which follows from the Lipschitz condition).
Further for global convergence we will assume Lipschitz continuity of the discretized energy functional. Note that since the metric varies smoothly on the manifold, then the discretized energy functional is also smooth and hence Lipschitz. Thus, all Finsler and Riemannian manifolds satisfies the above conditions, but an even larger class of energy functionals that does not necessarily stem from a Riemannian or Finsler metric will satisfy the requirements.
A.2 Proof of necessary conditions
Proposition 6.
The necessary conditions for a minimum in Eq. 6 is
where $\mu_{t}\in\mathbb{R}^{d}$ for $t=0,\dots,T-1$.
Proof.
The Hamiltonian of Eq. 6 is
$H_{t}(x_{t},u_{t},\mu_{t})=u_{t}^{\top}G(x_{t})u_{t}+\mu_{t}^{\top}(x_{t}+u_{t}).$ (17)
A.3 Proof of update scheme for Riemannian geometry
Proposition 7.
The update scheme in eq. 7 and eq. 9 for $u_{t},\mu_{t}$ and $x_{t}$ can in closed-form be written as
$\begin{split}&\mu_{T-1}=\left(\sum_{t=0}^{T-1}G_{t}^{-1}\right)^{-1}\left(2(a-b)-\sum_{t=0}^{T-1}G_{t}^{-1}\sum_{t>j}^{T-1}\nu_{j}\right),\\ &u_{t}=-\frac{1}{2}G_{t}^{-1}\left(\mu_{T-1}+\sum_{j>t}^{T-1}\nu_{j}\right),\quad t=0,\dots,T-1,\\ &x_{t+1}=x_{t}+u_{t},\quad t=0,\dots,T-1,\\ &x_{0}=a.\end{split}$
Proof.
From the update formula $\nu_{t}+\mu_{t}=\mu_{t-1}$ for $t=1,\dots,T-1$ we deduce that $\mu_{t}=\mu_{T-1}+\sum_{j>t}^{T-1}\nu_{j}$. From the update scheme $02G_{t}u_{t}+\mu_{t}=$ 0 we deduce that $u_{t}=-\frac{1}{2}G_{t}^{-1}\mu_{t}$, since $G_{t}$ is positive definite and hence has a well-defined inverse. Noting that $\sum_{t=0}^{T-1}u_{t}=b-a$ we can re-write the iterative formulas as
which proves the result. ∎
A.4 Proof of update scheme for Finslerian geometry
Consider the following control problem in the Finslerian case.
A.5 Proof of global convergence
To prove global convergence for the algorithm GEORCE we first show existence and uniqueness by proving two lemmas. Secondly, we prove that in each iteration the discretized energy functional is decreasing, unless a (local) minimum has been found. We will assume Lipschitz continuity of the metric tensor as described in 5.
[…]
@@ A.5 Proof of global convergence
Assume that $(x^{(i)},u^{(i)})$ is a feasible solution, then
@@
This shows that the linear combination in the state is feasible as it produces feasible state variables in terms of start and end point. Furthermore, each state variable is feasible, if they were determined from the linear combination of the control vectors, which proves that the linear combinations of the state and control variables are also feasible solutions. ∎
@@
$\begin{split}{\left.\kern-1.2pt\nabla_{x_{t}}E(x,u)\vphantom{\big{|}}\right|_{(x,u)=\left(x_{t}^{(i)},u_{t}^{(i)}\right)}}&=\mu_{t-1}-\mu_{t},\quad t=1,\dots,T-1,\\ {\left.\kern-1.2pt\nabla_{u_{t}}E(x_{t},u_{t})\vphantom{\big{|}}\right|_{(x_{t},u_{t})=\left(x_{t}^{(i)},u_{t}^{(i+1)}\right)}}&=-\mu_{t},\quad t=0,\dots,T-1.\end{split}$ (22)
@@
Since $G\left(x_{t}^{(i)}\right)$ is positive definite, the first summation is negative assuming that at least one non-zero vector $\{\Delta u_{t}\}_{t=0}^{T-1}$, which is the case, since we have assumed that $\left(x_{t}^{(i)},u_{t}^{(i)}\right)$ is not a (local) minimum. Now scale all vectors, $\left\{\Delta u_{t}\right\}_{t=0}^{T-1}$ with a scalar $10<\alpha<$ 1. Note that the term, $-2G\left(x_{t}^{(i)}\right)\Delta u_{t}$, is unaffected by this, since it equals $\mu_{t}+\nabla_{u_{t}}E(x,u)\left(x_{t}^{(i)},u_{t}^{(i)}\right)$, which is independent of $\alpha$. The first order Taylor order approximation is
@@
The discretized energy functional is a positive function, i.e. a lower bound is at least 0. Assume that the series $\left\{E^{(i)}\right\}_{i>0}$ is not converging to a (local) minimum. From lemma 9 it follows that $\left\{E^{(i)}\right\}_{i>0}$ is decreasing for increasing $i$, and since $E^{(i)}$ has a lower bound, the series $\left\{E^{(i)}\right\}_{i}$ can not be non-converging. Assume now that the convergence value for $\left\{E^{(i)}\right\}_{i}$, denoted $\hat{E}$, is not a (local) minimum, and denote the convergence point $(\hat{x},\hat{u})$. According to lemma 9 GEORCE will from the solution $(\hat{x},\hat{u})$ in the following iteration produce a new point $(\tilde{x},\tilde{u})$ such that $E(\tilde{x},\tilde{u})<\hat{E}$, which is contradiction as the series $\left\{E^{(i)}\right\}_{i}$ is assumed to be converging to $\hat{E}$. The series, $\left\{E^{(i)}\right\}_{i}$, will therefore converge, and the convergence point is a (local) minimum. ∎
Under the assumption of a strongly unique minimum point in Appendix A.1 GEORCE will converge to a local minimum by 3. Thus, it is not possible that GEORCE will ”jump” between local minimas.
A.6 Proof of local quadratic convergence
In this section we show that GEORCE has locally quadratic convergence to a (local) minimum under certain assumptions satisfied by the discretized energy functional. We will assume the assumptions in 5 hold. We first prove the following result, before proving quadratic local convergence.
@@ A.6 Proof of local quadratic convergence
Assume $f(x)$ is a smooth function. Let $z^{*}$ be a strongly unique (local) minimum point for $f$. Then
@@
Since $f(x)$ is a smooth function, then the first order Taylor approximation for $f(x)$ exists. Consider a point $z\in B_{\epsilon}(z^{*})\setminus\{z^{*}\}$, then
$f(z^{*})-f(z)=\langle\nabla f(z),z^{*}-z\rangle+\mathcal{O}\left(z^{*}-z\right)\left\lVert z^{*}-z\right\rVert.$
Since $z^{*}$ is a strongly unique (local) minimum point, then
$f(z^{*})-f(z)\leq-K\left\lVert z^{*}-z\right\rVert<0,\quad K>0,$
@@
$\frac{f(z^{*})-f(z)}{\left\lVert z^{*}-z\right\rVert}=\left\langle\nabla f(z),\frac{z^{*}-z}{\left\lVert z^{*}-z\right\rVert}\right\rangle+\mathcal{O}\left(z^{*}-z\right)\leq-K.$
Since $\lim_{z\rightarrow z^{*}}\mathcal{O}(z^{*}-z)=0$, then there exists $\epsilon>0$ such that
$\left\langle\nabla f(z),\frac{z^{*}-z}{\left\lVert z^{*}-z\right\rVert}\right\rangle<0,$
which implies that $\langle\nabla f(z),z^{*}-z\rangle<0$ for some $\epsilon>0$ for all $z\in B_{\epsilon}\left(z^{*}\right)\setminus\left\{z^{*}\right\}$. ∎
@@
If the discretized energy functional has a strongly unique (local) minimum point $z^{*}$ and locally $\alpha^{*}=1$, i.e. no line search, then GEORCE has locally quadratic convergence, i.e.

### v2

@@ 3.1 Deriviaton of GEORCE
GEORCE has complexity $\mathcal{O}\left(Td^{3}\right)$ due to matrix inversions along the discretized curve but scales only linearly in $T$, unlike e.g. quasi-Newton methods (Table 1). Although the computational complexity in the manifold dimension is of a lower order for quasi-Newton methods and gradient descent, GEORCE exhibits faster convergence. Since the Hessian of the discretized energy functional in Eq. 5 is sparse, the update step in the Newton method can be simplified in complexity compared to the standard Newton method that has complexity $\mathcal{O}\left(T^{3}d^{3}\right)$ (see Appendix B.2). However, the sparse Newton step requires second-order derivatives, unlike GEORCE, which only requires first-order derivatives of the discretized energy. Note that in Table 1, we assume that the iterations for the different methods remain well-defined in the chosen chart.
3.2 Convergence results
In this section, we show that GEORCE exhibits global convergence, i.e., converges to a local minimum, and that GEORCE exhibits local quadratic convergence under certain regularity of the discretized energy functional. To simplify the notation, we let $\langle\cdot,\cdot\rangle$, $\left\lVert\cdot\right\rVert$ and $\nabla$ denote the Euclidean inner product, norm and gradient in $\mathbb{R}^{d}$, respectively.
Global convergence
We will prove that GEORCE converges to a (local) minimum. To prove this, we will use the fact that the discretized energy functional is smooth and hence locally Lipschitz. This implies that the first order Taylor approximation of the discretized energy functional can be written as
$\Delta E=\langle\nabla E(z_{0}),\Delta z\rangle+\mathcal{O}\left(\Delta z\right)\left\lVert\Delta z\right\rVert,$ (12)
Thus, the term $\mathcal{O}\left(\Delta z\right)\left\lVert\Delta z\right\rVert$ can be re-written as $\mathcal{O}\left(\left\lVert\Delta z\right\rVert^{2}\right)$ locally.
To prove global convergence for the algorithm GEORCE we first show existence and uniqueness by proving two lemmas. Secondly, we prove that in each iteration the discretized energy functional is decreasing, unless a (local) minimum has been found.
For notation set $x^{(i)}=(a,x_{1}^{(i)},\dots,x_{T-1}^{(i)},b)$ and $u^{(i)}=(u_{0}^{(i)},u_{1}^{(i)},\dots,u_{T-2}^{(i)},u_{T-1}^{(i)})$ as the solution for iteration $i$ such that $(x^{(i+1)},u^{(i+1)})$ is the solution using the update scheme in proposition 2.
@@ 3.2 Convergence results
Assume that $\left(x^{(i)},u^{(i)}\right)$ is a feasible solution, then
@@
Initially, assume that $\left(x^{(i+1)},u^{(i+1)}\right)$ is well-defined in the local chart.
@@
This shows that the linear combination in the state is feasible as it produces feasible state variables in terms of start and end point. Furthermore, each state variable is feasible, if they were determined from the linear combination of the control vectors, which proves that the linear combinations of the state and control variables are also feasible solutions.
If $\left(x^{(i+1)},u^{(i+1)}\right)$ is not well-defined in the local chart, then the line-search in GEORCE will be able to determine a point belonging to the manifold and that will be on the line between $\left(x^{(i+1)},u^{(i+1)}\right)$ and $\left(x^{(i)},u^{(i)}\right)$ as the local chart is defined on an open set and $x^{(i)}$ is well-defined in the local chart. The new solution is then a feasible solution following the argument above. ∎
@@
$\begin{split}{\left.\kern-1.2pt\nabla_{x_{t}}E(x,u)\vphantom{\big|}\right|_{(x,u)=\left(x_{t}^{(i)},u_{t}^{(i)}\right)}}&=\mu_{t-1}-\mu_{t},\quad t=1,\dots,T-1,\\ {\left.\kern-1.2pt\nabla_{u_{t}}E(x,u)\vphantom{\big|}\right|_{(x_{t},u_{t})=\left(x_{t}^{(i)},u_{t}^{(i+1)}\right)}}&=-\mu_{t},\quad t=0,\dots,T-1.\end{split}$ (13)
@@
Since $G\left(x_{t}^{(i)}\right)$ is positive definite, the first summation is negative assuming that at least one non-zero vector $\{\Delta u_{t}\}_{t=0}^{T-1}$, which is the case, since we have assumed that $\left(x_{t}^{(i)},u_{t}^{(i)}\right)$ is not a (local) minimum. Now scale all vectors, $\left\{\Delta u_{t}\right\}_{t=0}^{T-1}$ with a scalar $0<\alpha<1$. Note that the term, $2G\left(x_{t}^{(i)}\right)\Delta u_{t}$, is unaffected by this, since it equals $\mu_{t}+\nabla_{u_{t}}E(x,u)\left(x_{t}^{(i)},u_{t}^{(i)}\right)$, which is independent of $\alpha$. The first order Taylor order approximation is
@@
[…]
It is not possible for GEORCE to jump between two local minimum points. To see this, assume that there are two local minimum points that have the same discretized energy, and assume that GEORCE is jumping between the two points, i.e., there is convergence to a local minimum in terms of the discretized energy functional according to Proposition 5, but not to a single local minimum point. In the GEORCE iteration, the line search will try to improve the discretized energy functional by finding a point closer to one of the two minimum points. Either this leads to a new point with improved discretized energy, or GEORCE stops concluding that a local minimum point has been found since no improvements in discretized energy were possible by line-search in accordance with Lemma 4. But this contradicts the assumption that GEORCE jumps between two minimum points with the same discretized energy. Thus, GEORCE converges to only one local minimum point.
Local convergence
We will show that GEORCE under certain regularity of the discretized energy functional exhibits local convergence in the sense that
$\exists\epsilon>0:\quad\exists c>0:\quad\forall z^{(i)}\in B_{\epsilon}(z^{*}):\quad\left\lVert z^{(i+1)}-z^{*}\right\rVert\leq c\left\lVert z^{(i)}-z^{*}\right\rVert^{2},$ (16)
for some $c>0$. We will assume the following regularity of the discretized energy functional
Assumption 6 (Local convergence).
  • We assume that the discretized energy functional $E(z)$ is locally strictly convex in the (local) minimum point $z^{*}=\left(x^{*},u^{*}\right)$ in the sense that
    $\exists\epsilon>0:\,\forall z\in B_{\epsilon}\left(z^{*}\right),z\neq z^{*}:\,\forall\alpha]0,1[:\,E\left(\left(1-\alpha\right)z+\alpha z^{*}\right)<\left(1-\alpha\right)E(z)+\alpha E\left(z^{*}\right),$
    where $B_{\epsilon}=\left\{z\,|\,\left\lVert z-z^{*}\right\rVert<\epsilon\right\}$.
  • Assume that the discretized energy functional $E(z)$ is smooth, and consider the first order Taylor approximation of the discretized energy functional
    Then, the term $\mathcal{O}\left(\Delta z\right)\left\lVert\Delta z\right\rVert$ can be re-written as $\mathcal{O}\left(\left\lVert\Delta z\right\rVert^{2}\right)$ locally.
Figure 3: Illustration of the convergence of GEORCE and baseline algorithms. We plot $\log\left\lVert z^{(i)}-z^{*}\right\rVert$ over $\log\left\lVert z^{(i+1)}-z^{*}\right\rVert$ for each iteration $i$ for the different methods. We use the convergence point of GEORCE as $z^{*}$. If the norms are 0, we set the value to 20-20. All initial curves are set to the same random deformation of a straight line similar to Appendix C.
Intuitively, it is expected that GEORCE exhibits local quadratic convergence. GEORCE solves iteratively quadratic minimization problems by fixing the metric matrix function. Since GEORCE exhibits global convergence, the metric matrix function in each iteration of GEORCE will converge to the metric matrix function at the local minimum. Close to the local minimum, the metric matrix function will be approximately constant, since it is smooth and hence continuous. At that point, GEORCE is simply solving a quadratic optimization problem, which is solved in one iteration similar to the Newton method, i.e,. has quadratic convergence. It is therefore expected that when GEORCE is sufficiently close to the minimum a large “dip” will occur. This is especially clear for the Euclidean space, where $G=I$ is independent of $x$. Therefore, GEORCE will find the exact solution after one iteration, as it simply solves a quadratic minimization problem. We illustrate the convergence of GEORCE and the baseline algorithms for $\mathbb{S}^{2}$ and $\mathbb{R}^{2}$ in Fig. 3. We first prove the following result, before proving quadratic local convergence.
@@
Assume $f(x)$ is a smooth function. Let $z^{*}$ be a (local) minimum point for $f$, and assume that $f$ is locally strictly convex in $z^{*}$. Then
@@
Since $f$ is locally strictly convex in $z^{*}$, it follows that
$\exists\epsilon>0:\quad\forall z\in B_{\epsilon}\left(z^{*}\right),z\neq z^{*}:\quad\forall\alpha\in]0,1[:\quad f\left(\left(1-\alpha\right)z+\alpha z^{*}\right)<\left(1-\alpha\right)f(z)+\alpha f\left(z^{*}\right),$
@@
$\exists\epsilon>0:\quad\forall z\in B_{\epsilon}\left(z^{*}\right):\quad\forall\alpha\in]0,1[:\quad\frac{f\left(z+\alpha\left(z^{*}-z\right)\right)-f(z)}{\alpha}<f\left(z^{*}\right)-f(z).$
As $f$ is locally strictly convex in the minimum point $z^{*}$
$\exists\epsilon>0:\quad\forall z\in B_{\epsilon}\left(z^{*}\right),z\neq z^{*}:\quad f\left(z^{*}\right)-f(z)<0,$
which follows from the fact that if there exists a $z^{\prime}\in B_{\epsilon}\left(z^{*}\right)$ with $f\left(z^{\prime}\right)=f\left(z^{*}\right)$, then since $f$ is locally strictly convex in $z^{*}$, then
$\forall\alpha\in]0,1[:\quad(1-\alpha)f\left(z^{\prime}\right)+\alpha f\left(z^{*}\right)=f\left(z^{*}\right)>f\left(\left(1-\alpha\right)z+\alpha z^{*}\right).$
However, this contradicts that $z^{\prime}$ and $z^{*}$ are local minimum points. Since $f$ is a smooth function, then the first order Taylor approximation gives
$\frac{f\left(z+\alpha\left(z^{*}-z\right)\right)-f(z)}{\alpha}=\left\langle\nabla f(z),z^{*}-z\right\rangle+\mathcal{O}\left(\alpha\left(z^{*}-z\right)\right)\left\lVert z^{*}-z\right\rVert<f\left(z^{*}\right)-f(z)<0.$
In the limit, then $\lim_{\alpha\searrow 0}\mathcal{O}\left(\alpha\left(z^{*}-z\right)\right)=0$, which implies that $\left\langle\nabla f(z),z^{*}-z\right\rangle<0$. ∎
@@
If the discretized energy functional is locally strictly convex in the (local) minimum point $z^{*}$ and locally $\alpha^{*}=1$, i.e. no line search, then GEORCE has locally quadratic convergence, i.e.